| 
Group Selection Page
HELP
Controls Group
ITDM
Vacuum
WP78
FXE
SPB
HED
SQS
SXP
PSPO
XO
TS
Migrated2Zulip
Archived2GitlabPages
Archived2PDF
MigratedByGroup
 | 
| 
SCS
Photon Commissioning Team
XRO
DOC
WP74
MID
XO_mig
XO_Operation
SQS Migrated2Zulip
DET
Sample Environment
 | 
| 
Technical integration and coordination
 | 
|   Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  | 
Not logged in | 
  | 
   | 
| 
Message ID: 43
    Entry time: 15 May 2019, 12:53
 |  
| Author: | 
Adriano Violante  | 
 | Location: | 
General  | 
 | Severity: | 
Information  | 
 | Subject: | 
Minutes of the Technical integration meeting 15.05.19   | 
 
  |  
 
 | 
SPB/SFX 
	- Nano KB mirrors and CSS controls are OK
 
	- Digitizers have been requested. HW needs to be installed and set up and the timeline has to be clarified betweek BF and SPB. Not possible to have them running on short notice.
 
	- 
	
Comment from AS: Trigger Test with pulse on demand to test the conditional trigger (5Hz like operation) is done, but failed to deliver a real conditional trigger. The Karabo device is adapted to switch between the present mode and conditional trigger from Karabo (although some issue were identified). The X2Timer configuration still need adjustment, BF is on it. 
	 
 
FXE 
	- Re shuffling of PLC CPUs drawers will occur during the next winter shutdown.
 
 
SQS 
	- MPODS power supplies are behaving strangely. CY will have a look.
 
	- pnCCD: Commissioning should start today. Maxi cage for interlocks will be installed today by SQS.
 
 
SCS 
 MID 
	- Axes for users operation being commissed
 
	- Communication between VAC and AE concerning the electronics for the cryo coolers - heaters has to be improved since AE was not involved during the last ZZ intervention on monday. MD will discuss later today with NJ.
 
 
HED 
	- Loop 6 can be operated in its completeness by middle of next week.
 
	- Laser beam has to be deivered to SCS in 2 weeks. Fw has to be uploaded in the PLCs. HW is ready, commissioning has to start.
 
	- There is the request to make the timing signal independent between the SASEs. Will be discussed offline with AE. It should be possible in the winter shutdown. ==> BF is talking to the involved people.
 
 
XRO 
XPD 
VAC 
DET 
	- Karabo 2.4.1. improves the pipieline stability. The deployment is strongly suggested to MID and SPB. ==> Pipeline for SPB has been deployed. Minor issues being addressed. MID has to folllow.
 
 
AE 
ITDM 
EET 
	- Schedule for the Loop 10 and 11 of HED cables installation is confirmed. IN KW25 would be possible to bring the loop up.
 
 
CAS 
SRP 
TS 
AoB  |